OpenStudy (jkbo):

Please help.What is the solution to the system of equations? 2x - 3y -2z=-7 x-3y-3z=4 x-2y-z=-3 a.(–8, –1, 3) b.(–8, –1, –3) c.(–8, 1, –3) d.(8, –1, –3)

OpenStudy (texaschic101):

elimination method ....are you familiar with this ?

OpenStudy (texaschic101):

2x - 3y - 2z = -7 x - 3y - 3z = 4 -->(-2) --------------- 2x - 3y - 2z = -7 -2x + 6y + 6z = -8 (result of multiplying by -2) ---------------add 0 + 3y + 4z = - 15 3y + 4z = -15 x - 3y - 3z = 4 -->(-1) x - 2y - z = -3 ----------------- -x + 3y + 3z = -4 (result of multiplying by -1) x - 2y - z = -3 -----------------add 0 + y + 2z = -7 y + 2z = -7 3y + 4z = -15 y + 2z = -7 -->(-2) ---------------- 3y + 4z = -15 -2y - 4z = 14 (result of multiplying by -2) ----------------add y = -1 now sub -1 in for y y + 2z = -7 -1 + 2z = -7 2z = -7 + 1 2z = -6 z = -3 now sub known values into one of the original equations 2x - 3y - 2z = -7 2x - 3(-1) - 2(-3) = -7 2x + 3 + 6 = -7 2x + 9 = -7 2x = -9 -7 2x = - 16 x = -8 lets check our answers... x - 3y - 3z = 4 -8 - 3(-1) - 3(-3) = 4 -8 + 3 + 9 = 4 -8 + 12 = 4 4 = 4 (correct) answer : x = -8, y = -1, z = -3 or (-8,-1,-3)

OpenStudy (jkbo):

Thank you

OpenStudy (texaschic101):

your welcome :)

OpenStudy (anonymous):

make two equations by eliminating any variable say x from three equations, taken two at a time.then solving for y and z from the two equations. finally substitute in any one of the original three equations and get the value of x

OpenStudy (texaschic101):

you do have to be real careful with elimination because it is easy to make mistakes.